OpenStudy (anonymous):

If f(x) = 1/3x − 2 and g(x) = −2x − 3, find (f • g)(6)

OpenStudy (anonymous):

\[f ( g(x)) = f( -2x - 3) = (1/3)(-2x - 3) - 2\] \[= (-2/3)x - 1-2\] \[= (-2/3)x - 3\] You can plug any value of x into this equation, so in this instance plug in x = 6 \[= [(-2/3)\times6] - 3 \] \[= -7\]